OpenStudy (anonymous):

(2-square root of 7)(2+square root of 7)

OpenStudy (angela210793):

(a-b)*(a+b)=a^2-b^2 try to solve it now pls

OpenStudy (angela210793):

=2^2-(sqr7)^2=4-7=-3